(1) Suspension health check for IX, EX, PR3, PR2 (IMMT1/2) : (Ushiba-kun and Takahashi-san)


A01: Cable check: Type-A:by R.Takahashi-san), Sato-san has raw data -> OK 空欄部分の意味は?線が切れてるかつながってない状態 → 正常なことを明記お願いします。

Cryo-payload by Ushiba 内外のケーブルチェック:OK FeedThroughのところで、10%以上のばらつき有り。 TBDはスキップ。モデルがないので。

A02 Porality Check Type-A/C-Payload: OK Eddy Current Damping もOK

A03 RubbingCheck Type-A: 〇の意味は?mmでわかりますか?→ 全開になってるはず。その時の間隔は?設計上は?F0~ cm orderのはず。最低でもcmあるはず。写真は?ないです。図を入れてわかるようします。 F1~F3は見えないのでEY以外はデータはない。

C-Payload: 目視確認、TM bottom1~4 はせめぎみ。冷やしても広がるセンス。 MN関係は 3.5mm 開けすぎると磁石が当たるので、それを避けるため それ以外は10mm程度 ギャップに関する理由づけもドキュメントにお願いします。 たたいてみて、触れてないこともチェック。

A0:4Range and Calibration Type-A: Fishing rod dynamic raange and graphs。軸の単位もお願いします(横軸 ステップ数、縦軸 カウント) LVDT calibrationは金尺以外の方法で、精度を上げられませんか? LVDTに関し、動かせる範囲が明示。Keystoneの高さをCalibration後のLVDTの値から読み取った値(単位mm) Catuator Rangeについては、計測すみで、書き込んでないだけ。効率は伝達関数から見積もるつもりなので、あとで計算。

C-Payload moving mass Range 確認:IXしかできてない。どっちの方向も1mmは動かせる。実験室では、1mmradは動かせていた。再現性はそんなにない。 これで十分なはず。EXではまだやってない。今回は、確実に動くことを目指した。冗長性のため2台入れている。 Actuator Efficiency: 一個を代表して測定

A05: Watchdog Setting 後でもできるのでSkip

A06:Height Measurement type-A: LVDTカウント値のセットポイント。klog参照できるように書き込む。

A08: Calibration Type-A: LVDT:根拠のリンクをはってください。 Cryo-P: Photo センサーは不要 Oplev Yokozawa君のoplev summary documentation 参照でOKのはず。

A09: Spectrum No strange. So OK.

A10: Transfer Function No strange. So OK.

IX -> OK Couplingは少し変化してるが、あとで、decoupleする。 Couplingが大きいものはピックアップした。 BFY 20 -> 30 mHz is as predicted.

EX Couplingが増えたものが多いが、、、あとで直す。 F1 damperの高さを変えると、劇的にCouplingが変わるのはIXで起こっているので、そのせいかも。問題ないであろう。 BF LVDTの値も変わっているので、センシング効率が変か?(Fig11)

A11 TF in the alignment state. 真空引きの後でいいのでは?

LVDTについて 測定忘れは無くて、測定値については?ばらつきはたくさんある。今のところ問題は起こってない。 特性が同じであるべきなのはその通りがもう戻れない。


(2)Oplev Check for IX, EX, PR3, PR2 (IMMT1/2) (Yokozawa-kun)

Beam path clearance: well away from obstacle structures? The beams hit around in the center of steering mirrors?

QPD centering: Beams are set around the QPD center? (This centering will satisfy the sensors dynamic ranges in the suspension systems? More offload is necessary?)

Clean signals: No seriously large line noises / and white noises?

Covering: the covers kill light from outside well ?

Beam source degradation: No serious power/beam quality degradation until now?

(3)CAL (ChenDan-kun and his selected person)

Please explain what you did after the review.

Beam diameter checks.

Beam path realignment after the accident,

Spectrum health check,

Mirror actuators health check,

(4)Tcam (Yokozawa-kun and ChenDan kun, and Yamamoto-kun?)

T-cam system stability,

Image can be obtained without illuminators?

Focus and inclination are proper?

(5)TMSX optical system (Akutsu-kun and his selected persons)

Trans image are captured and its image is clear inside the image area?

PD was set to show the max output when we center the green beam on the ETMY mirror?

Clear optical paths on the way to the TMSX detectors(PD/QPDs)?

(6)Baffle PDs (Yokozawa-kun and his selected persons)

Outputs are alive, except for one dead one?

Clear signals without any serious line noises and high noise floors?

(7)Baffles (Akutsu-kun and his selected persons)

Setting positions: designed positions?

Freely suspended?: position sensors are working properly? (spectrum is OK?)

The resonance damping level is proper?: The proper damping should be realized when the systems are cooled.

temp data: The temp sensors work well?

(8)Vacuum system (Kimura-san / Uchiyama-kun and their selected persons)

GV(BS-IXC) is closed?

The vacuum sensors are working well?

The flanges that used to be opened are properly closed? Do you have check lists?

The remedy point check between EXC and TMSX

Not to forget to put level gauges to sense the V-tank height

check of oplev windows, especially.

(9)Vacuum tank cleaning/ lost items

IXA/EXA: chendan-kun, yokozawa-kun

IEXC: Cryo members.

lost items check: miyoki and so on.

(10)Cryo-system (Kimura-san and his selected persons)

Temp data are properly obtained?

Insulators are attached properly?

The new cryocoolers are ever tested even for very short duration in the past after setting?

Heaters for windows on the cryo-shields properly works?

(11)Optics and stray light check in POP/REFL/IMCREFL/MCE tables (Michimura-kun/Akutsu-kun and his selected members?) Maybe this work cannot be finished before 10th, but in 13th week in December, if possible.

Used optics are well understood about their optical properties and damage thresholds?

Stray lights are well dumped?

KAGRA/Upgrades/2021XarmClosing (last edited 2021-12-10 14:17:04 by miyoki)